OpenStudy (anonymous):

Your class collected more than 250 cans of food in two weeks for an annual food drive in the first week 115 cans of food was collected how many cans c of food were collected in the second week write an solve the inequality

OpenStudy (anonymous):

115 + c < 250 c < 135 115 + c > 250 c > 135 c - 115 > 250 c > 365 c + 115 > 250 c > 365

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@Er.Mohd.AMIR

OpenStudy (anonymous):

@dr0zier99 @PuppyLover784

OpenStudy (anonymous):

@Er.Mohd.AMIR help me out please

OpenStudy (er.mohd.amir):

B

OpenStudy (anonymous):

u sure

OpenStudy (er.mohd.amir):

115+c>250 c>250-115 c>135

OpenStudy (er.mohd.amir):

or we can say c>250 c>250-115 c>135

OpenStudy (anonymous):

can u help me with two more please please ok

OpenStudy (er.mohd.amir):

ok

OpenStudy (anonymous):

A custom designer made 10 customs that each used 5.8 yards of fabric if she had 238.2 yard of fabric to start with about how much fabric does the custom designer have left with? write an equation and estimate the amount of fabric she has left.

OpenStudy (anonymous):

@Loser66

OpenStudy (er.mohd.amir):

LET the left is x then 10*5.8+x=238.2

OpenStudy (er.mohd.amir):

x=180.2

OpenStudy (anonymous):

so 10*5.8+x=238.2 is the equation and x=180.2 is the answer right

OpenStudy (er.mohd.amir):

yes

OpenStudy (er.mohd.amir):

58+x=238.2 x=238.2-58 x=180.2